OpenStudy (anonymous):

(3n+5)(2n-6)=0 Is this FOIL? How do I solve for n?

hero (hero):

Use zero product property ab = 0 to solve this

hero (hero):

If the product of two numbers, a and b, is zero, then a = 0, or b = 0

hero (hero):

so 3n + 5 = 0 or 2n - 6 = 0

OpenStudy (anonymous):

either 3n+5 = 0 or 2n-6 = 0 solve these two equations and you've got it

OpenStudy (anonymous):

FOIL is a memory aid for expanding the function - best to use distributive rule - multiply second bracket by first term and then by secoind and simplify